SUBMILLIMETER SPECTRUM OF PYRIDINE USING FASSST

F. L. BETTENS, Department of Chemistry, Australian National University, Canberra ACT 0200, Australia; R. P. A. BETTENS, Research School of Chemistry, Australian National University, Canberra ACT 0200, Australia; \underline{S. ALBERT} and F. C. DE LUCIA, Department of Physics, The Ohio State University, 174 West 18th Ave., Columbus, OH 43210-1106.

With the recent development of the Fast Scan Submillimeter Spectroscopic Technique (FASSST), the acquisition of submm-wave spectra over the 260 - 360 GHz region in a matter of seconds has become routine. This technique has so far been utilised to take rotational spectra of ethylene oxide and methyl mercaptan. Here we report the measurement of the rotational spectrum of pyridine. The newly taken spectrum has been combined with previous laboratory data and analysed. Due to the very large number of transitions measured, a partially automated technique has been developed to speed and aid in the assignment and fitting.